You are reading a news article discussing bank-fraud victimization. The report features the story of one victim who was 20 years old when she discovered her bank account had been compromised and her money had been stolen. Because you are familiar with the National Crime Victimization Survey's Identity Theft Supplement, you know that the mean age of bank-fraud victims is 45,52 (s= 15,26). The standard deviation is 15,26. Based on this, is the victim in the report typical of bank-fraud victims in terms of age? How do you know? Do you trust that her experiences are representative of other-bank-fraud victims? Explain your answer
Answer #1

Please don't hesitate to give a "thumbs up" in case you're satisfied with the answer

Lets check the Z score

Z = (Xbar - Mu)/stdev

= (20-45.52)/15.26

= -1.67

p-value = P(Z<-1.67) = .047

Since this is less than .05 ( usual cutoff taken for an event to be unusual), we can say that this report is NOT a typical one, her experiences are not a representative of other - bank- fraud victims.